# The $\alpha$–regression for compositional data: a unified framework for standard, spatially-lagged, spatial autoregressive and geographically-weighted regression models

### v3

@@ 3.1.4 Properties of the SMEs
    $\sum_{i=1}^{D}\frac{\partial\mu_{i}}{\partial x_{k}}=0\quad\text{and}\quad\sum_{i=1}^{D}\frac{\partial\mu_{i}}{\partial(\mathbf{W}\mathbf{x})_{k}}=0$ (22)
@@ 3.2.2 SMEs
$\frac{\partial\mu_{i\ell}}{\partial x_{ik}}=\begin{cases}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\cdot[S(\rho)^{-1}]_{ii}&\text{for }\ell=1\\[8.61108pt] \mu_{i\ell}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]\cdot[S(\rho)^{-1}]_{ii}&\text{for }\ell=2,\ldots,D.\end{cases}$ (25)
@@
$\frac{\partial\mu_{i\ell}}{\partial x_{jk}}=\begin{cases}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}&\text{for }\ell=1\\[8.61108pt] \mu_{i\ell}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}&\text{for }\ell=2,\ldots,D\end{cases}$
@@ 3.2.3 Prediction of new values
Denote the new $m$ covariate values by $\mathbf{X}^{new}$ located at new, unseen in the model, coordinates. We stack the the new covariate values under the observed ones to create the augmented design matrix
@@ 3.3.4 SMEs
Jacobian vector at location $(\nu_{i},u_{i})$
where $r\in\{1,\ldots,d\}$ and $s\in\{1,\ldots,p\}$.
@@ 4.1 Agricultural economics data set
Figure LABEL:fig_real(a) shows the location of Thessaly region in Greece, and Figure LABEL:fig_real(b) shows the locations of the farms. The majority of the farms cultivate cereals and only few farms hold grapes and wine. Specifically, 84.52% of the farms cultivate cereals, 50.00% cultivate Cotton, 40.48% maintain tree crops, 81.55% hold other annual crops and pasture, and finally only 16.67% of the farms own grapes and wine.
@@ 4.2 Meuse river data set
We then fitted the regression models using the optimal parameters obtained based on the CV protocol and computed the correlations (component-wise) between the observed and fitted compositions. Table 2 contains these correlations. Evidently, the $\alpha$–SAR model seems to have achieved the best fit. It is important to highlight that the spatial autoregressive parameter $\rho$ of the $\alpha$–SAR model was equal to -0.206 with a standard error equal to 0.164. Table 4 contains the correlations between each pair of the observed and fitted components for each of the four regression models. The $\alpha$–regression, $\alpha$–SLX and the $\alpha$–SAR model have performed equally well.

### v4

@@ 3.1.4 Properties of the SMEs
    $\sum_{i=1}^{D}\frac{\partial\mu_{i}}{\partial x_{k}}=0\quad\text{and}\quad\sum_{i=1}^{D}\frac{\partial\mu_{i}}{\partial(\mathbf{W}\mathbf{x})_{k}}=0$
@@ 3.2.2 SMEs
$DSME_{ilk}=\frac{\partial\mu_{i\ell}}{\partial x_{ik}}=\begin{cases}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\cdot[S(\rho)^{-1}]_{ii}&\text{for }\ell=1\\[8.61108pt] \mu_{i\ell}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]\cdot[S(\rho)^{-1}]_{ii}&\text{for }\ell=2,\ldots,D.\end{cases}$
@@
$IDSME_{iljk}=\frac{\partial\mu_{i\ell}}{\partial x_{jk}}=\begin{cases}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}&\text{for }\ell=1\\[8.61108pt] \mu_{i\ell}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}&\text{for }\ell=2,\ldots,D\end{cases}$
@@ 3.2.3 Prediction of new values
Denote the new $m$ covariate values by $\mathbf{X}^{new}$ located at new, unseen in the model, coordinates. We stack the the new covariate values under the observed ones to create the augmented design matrix (Goulard et al., 2017)
@@ 4.1 Agricultural economics data set
Figure 1(a) shows the location of Thessaly region in Greece, and Figure 1(b) shows the locations of the farms. The majority of the farms cultivate cereals and only few farms hold grapes and wine. Specifically, 84.52% of the farms cultivate cereals, 50.00% cultivate Cotton, 40.48% maintain tree crops, 81.55% hold other annual crops and pasture, and finally only 16.67% of the farms own grapes and wine.
Refer to caption Refer to caption
(a) Region of Thessaly within Greece. (b) The locations of the 168 farms.
Figure 1: The Thessaly region in Greece.
@@ 4.2 Meuse river data set
We then fitted the regression models using the optimal parameters obtained based on the CV protocol and computed the correlations (component-wise) between the observed and fitted compositions. Table 2 contains these correlations. Evidently, the $\alpha$–SAR model seems to have achieved the best fit. It is important to highlight that the spatial autoregressive parameter $\rho$ of the $\alpha$–SAR model was equal to -0.206 with a standard error equal to 0.164. Table 4 contains the correlations between each pair of the observed and fitted components for each of the four regression models. The $\alpha$–SLX regression model has outperformed the rest.
@@ Appendix G Standard error of the SMEs of the $\alpha$–SAR model
Jacobian vector for the SMEs of the GWαR model at location $(\nu_{i},u_{i})$
where $r\in\{1,\ldots,d\}$ and $s\in\{1,\ldots,p\}$.